# Trigonal warping enables linear optical spectroscopy in single-valley superconductors

### v1

@@ Abstract
In superconductors with multiple pairing channels, Bardasis-Schrieffer modes and clapping modes arise as fluctuations in channels whose angular momenta differ from that of the pair condensate. Crystal symmetries often impose selection rules which keep these modes optically dark. We show that if pairing occurs around a single Fermi surface, trigonal warping renders both of these modes, as well as the quasiparticle excitation gap, visible in the longitudinal and Hall optical responses. Our results suggest that rhombohedral graphene multilayers, which are believe to host the required ingredients, might offer an ideal setting for the study of exotic superconducting collective modes.
@@
When multiple pairing channels compete for the title of superconducting ground state, collective order-parameter modes generically emerge. For example, in an $s$-wave superconductor, subdominant attraction in a $d$-wave channel can bind an in-gap mode corresponding to $d$-wave pair fluctuations, known as a Bardasis-Schrieffer (BaSch) mode [1]. Similarly, when a superconductor breaks time-reversal symmetry ($\mathcal{T}$), fluctuations in the time-reversed partner to the condensed channel manifest as generalizations of the clapping mode found in superfluid 3He [2]. By signaling the presence of uncondensed pairing channels, collective modes carry valuable information about their host superconductor [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 14, 31, 32, 26, 33, 34, 35]. However, BaSch and clapping modes are often invisible to linear optical response in the long-wavelength limit $\mathbf{q}\rightarrow 0$ due to crystal symmetries  [^1] Refs. [4, 5, 6, 11] argued that linear optical response could detect clapping modes because the decay of an incident electromagnetic field at the boundary of a superconductor produces $\mathbf{q}$ components up to the inverse penetration depth. Bulk probes at $\mathbf{q}=0$ remain desirable., motivating techniques such as Raman [16, 17, 18, 19, 20, 21, 22, 23, 24, 25] or tunneling spectroscopy [26, 28, 29].
Previous studies have shown how linear electromagnetic coupling to superconducting collective modes can survive the long-wavelength limit in certain multi-band superconductors [37, 30, 14, 31, 32], for example, or when the gap function breaks in-plane rotational symmetries [13]. An alternative is to apply a supercurrent [38, 39, 40, 32, 41]: Uniform motion of the condensate enacts a Galilean transformation, breaking the antisymmetry between velocities of paired electrons, thereby coupling light to internal Cooper pair dynamics even at $\mathbf{q}=0$. In this Letter, we show how trigonal warping can enable a similar effect when pairing occurs around a single Fermi surface (see Fig. 1), rendering BaSch and clapping modes visible to any probe sensitive to the frequency-dependent superfluid stiffness. Such probes include longitudinal and Hall conductivity, polar Kerr effect [42], reflectometry [43], or circuit quantum electrodynamics [44, 45, 46]. Crucially, the optical couplings enabled by trigonal warping do not require any background field or supercurrent.
@@
Rhombohedral four-, five-, and six-layer graphene (RnnG) may provide precisely the ingredients required for the optical couplings we discuss. Recent experiments [47, 48, 49] have revealed several superconducting phases that emerge from spin-and-valley-polarized normal states, each hosting a single $C_{3}$-symmetric Fermi surface. Theory suggests that some of these superconductors are chiral [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. They might therefore host clapping modes (in addition to BaSch modes). We discuss the implications of our results for RnnG below, after introducing a simple model that captures our coupling mechanism.
[…]
@@
with attractive couplings $g_{1}\geq g_{2}>0$, and where the factor of $1/2$ is for convenience. (Including more channels is conceptually straightforward.) $\delta g=g_{1}-g_{2}$ quantifies the competition between pairing channels at the critical temperature; $\delta g=0$ means perfect competition.
@@
The gap equation arises as the saddle-point condition for the $\Delta^{(\ell)}$ fields after integrating out the fermions. Our analysis holds when inversion breaking is weak enough that $|\tilde{\xi}_{\mathbf{k}}|<E_{0,\mathbf{k}}\equiv\sqrt{\xi_{0,\mathbf{k}}^{2}+|\Delta_{\mathbf{k}}|^{2}}$ at all relevant momenta, where $\Delta_{\mathbf{k}}$ is the mean-field gap function (i.e., the saddle-point value of $\Delta_{\mathbf{k},q=0}$). For a nodeless state with strong pairing interactions, this condition can be self-consistently satisfied everywhere in $\mathbf{k}$ space. In that case, surprisingly, $\tilde{\xi}$ has no impact on pairing and disappears from all calculations in the long-wavelength limit at $T\rightarrow 0$ [64], except for its contribution $\propto\tilde{\mathbf{v}}_{\mathbf{k}}\tau^{z}$ to the paramagnetic vertex (more on this below). We assume that the $\ell=1$ channel condenses, producing the pairing gap
@@
$\Delta^{(1)}(x)=e^{i\theta(x)}\left[\Delta_{\text{mf}}+h(x)\right]$ (7a)
and
@@
$h(x)$ and $\theta(x)\in\mathbb{R}$ are respectively the Higgs and Anderson-Bogoliubov-Goldstone modes. The interpretation of $a(x)$ and $b(x)\in\mathbb{R}$ depends on the pairing channels chosen. If $\chi^{(1)}_{\mathbf{k}}=\overline{\chi}^{(2)}_{\mathbf{k}}$, i.e., if the two pairing channels are time-reversed partners, then $a$ and $b$ are clapping modes. If instead $\chi^{(1)}_{\mathbf{k}}$ and $\chi^{(2)}_{\mathbf{k}}$ come from unrelated irreps, e.g. $p+ip$ and $f+if$ (or $s\text{-wave}$ and $d\text{-wave}$ in a spin-singlet model), then $a$ and $b$ would instead constitute BaSch modes. It will be useful to collect the fields $h,a$ and $b$ into $\Phi=\begin{pmatrix}h&a&b\end{pmatrix}^{T}$.
We now couple to electromagnetism through the standard minimal substitution. Electromagnetic linear response is controlled by the effective action at bilinear order in the gauge potential $A=(A^{0},\mathbf{A})$, so we drop all terms at third order or higher. We may remove the overall order-parameter phase $\theta(x)$ by a gauge transformation [65, 66, 67], which combines $A$ and the spacetime-gradient of $\theta$ into the gauge-invariant field $\mathcal{A}=(\mathcal{A}_{0},\boldsymbol{\mathcal{A}})\equiv(A_{0}+\frac{1}{2e}\partial_{\tau}\theta,\mathbf{A}-\frac{1}{2e}\gradient\theta)$. Note that $\theta$ will ultimately play no role in our calculation when we focus on $\mathbf{q}=0$ optical response [64]. We then reorganize the action for an expansion around mean field,
@@
Here $\tau^{0}$ and $\tau^{z}$ are the identity and third Pauli matrix in Nambu space. Equation (14) is analogous to supercurrent-enabled optical coupling [38, 39, 40, 32, 41]: $\tilde{\mathbf{v}}$ acts in this equation similarly to a background supercurrent at velocity $\mathbf{v}_{s}$, since both are inversion-even contributions to the electronic velocity which therefore contribute a $\tau^{z}$ to the paramagnetic vertex. Like a supercurrent, trigonal warping can make bright the otherwise-dark order-parameter modes $a$ and $b$. (Unlike a supercurrent [38, 39, 68, 41], trigonal warping does not optically activate the Higgs mode [64].) Performing the Matsubara integral at $T\rightarrow 0$, we find
@@
where $E_{0,\mathbf{k}}=\sqrt{\xi_{0,\mathbf{k}}^{2}+|\Delta_{\mathbf{k}}|^{2}}$, and the vector on the right-hand-side is indexed by $\phi=h,a,b$. Equation (15) assumes only that $|\tilde{\xi}|<E_{0}$ everywhere, and that $\Delta_{\mathbf{k}}=\Delta_{\text{mf}}\chi^{(1)}_{\mathbf{k}}$.
@@
$\tilde{\mathbf{v}}_{\mathbf{k}}=\partial_{\mathbf{k}}\tilde{\xi}_{\mathbf{k}}=\eta\mathbf{k}^{2}(\cos 2\varphi,-\sin 2\varphi).$ (17)
Consider the couplings in Eq. (15). The Higgs component vanishes since $\tilde{\mathbf{v}}$ averages to zero over the Fermi surface, and eveything else in the integrand has full rotational invariance. However, the integrands for the clapping mode components include the factor $\overline{\chi}^{(1)}\chi^{(2)}=e^{-2i\varphi}$ which balances the angular dependence of $\tilde{\mathbf{v}}$. Those modes become optically active:
@@
We plot the renormalized optical conductivity tensor,
@@
along with the longitudinal stiffness, $K^{xx}_{\mathbf{q}=0}$, in Fig. 1. The real part of the longitudinal conductivity consists of $\delta$-functions at DC ($\Omega=0$) and the clapping mode frequency $\Omega_{c}$, and a sharp peak at the quasiparticle gap $2\Delta_{\text{mf}}$. Breaking the degeneracy between pairing channels ($g_{1}>g_{2}$) pushes the clapping mode up in energy from $\sqrt{2}\Delta_{\text{mf}}$. (Within our model, the $a$ and $b$ modes are exactly degenerate for all values of $\delta g$, so they contribute a single pole; they may split when the gap function breaks rotational symmetry [13]). The Hall conductivity also exhibits sharp features at $\Omega_{c}$ and $2\Delta_{\text{mf}}$; a non-zero Hall response is permitted since the trigonally warped Fermi surface breaks inversion symmetry [69, 70, 71]. Inversion symmetry also typically prevents strong response from quasiparticle excitations across the gap in clean superconductors [72], but trigonal warping activates it along with the clapping modes, analogously to a supercurrent [73, 68].
@@
Application to rhombohedral graphene—We now discuss the implications of our results for RnnG. In the four, five and six layer cases, experiments have reported superconductors that condense from spin-and-valley-polarized normal states [47, 48, 49], suggesting that pairing occurs within a single flavor. Fermion anticommutation then forces the gap function to be either chiral or nodal. Theoretical models often prefer a chiral solution as it admits a full gap, which increases the condensation energy relative to a nodal state [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. Since the Fermi surface in each valley is trigonally warped, RnnG likely meets the criteria we have discussed for optical coupling to condensate modes.
@@
If a chiral superconductor spontaneously breaks $\mathcal{T}$ (i.e., if it arises from a $\mathcal{T}$-symmetric normal state), one generically expects a clapping mode comprising fluctuations in the time-reversed partner to the condensed channel. In superconductors descended from $\mathcal{T}$-broken normal states (like RnnG), the band-projected pairing interactions will also break $\mathcal{T}$, and attraction in the time-reversed channel could be much weaker than in the condensed channel. Another pairing channel (in a different irrep) could also have comparable attraction to the ground state [51]. As a result, a BaSch mode might arise at lower energy than the clapping mode.
@@
Identifying the character of any collective mode observed in experiment could help gain insight into the underlying pairing interactions. The BaSch modes we considered change pair angular momentum by two units while preserving its direction, while the $p+ip$ clapping mode reverses the pair angular momentum. Therefore, although in our calculations the two modes appear identically in the superfluid stiffness, we expect them to couple differently to an out-of-plane magnetic field. Future work should study this possibility within microscopic models.
@@
Our results extend the zoo of collective modes detectable in linear optical spectroscopy without background supercurrents. We argued that valley-polarized superconductors in RnnG are a promising platform to test our ideas. Future work will consider more realistic models of RnnG and investigate how experimental tuning parameters—such as charge density, perpendicular displacement fields, and magnetic fields—affect their collective mode spectra.

### v2

@@ Abstract
In superconductors with multiple pairing channels, Bardasis-Schrieffer modes and clapping modes arise as fluctuations in channels whose angular momenta differ from that of the pair condensate. Crystal symmetries often impose selection rules which keep these modes optically dark. We show that if pairing occurs around a single Fermi surface, trigonal warping renders both of these modes, as well as the quasiparticle excitation gap, visible in the longitudinal and Hall optical responses. Our results suggest that rhombohedral graphene multilayers, which are believed to host the required ingredients, might offer an ideal setting for the study of exotic superconducting collective modes.
@@
When multiple pairing channels compete for the title of superconducting ground state, collective order-parameter modes generically emerge. For example, in an $s$-wave superconductor, subdominant attraction in a $d$-wave channel can bind an in-gap mode corresponding to $d$-wave pair fluctuations, known as a Bardasis-Schrieffer (BaSch) mode [1]. Similarly, when a superconductor breaks time-reversal symmetry ($\mathcal{T}$), fluctuations in the time-reversed partner to the condensed channel manifest as generalizations of the clapping mode found in superfluid 3He [2]. By signaling the presence of uncondensed pairing channels, collective modes carry valuable information about their host superconductor [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 14, 31, 32, 26, 33, 34, 35, 36]. However, BaSch and clapping modes are often invisible to linear optical response in the long-wavelength limit $\mathbf{q}\rightarrow 0$ due to crystal symmetries  [^1] Refs. [4, 5, 6, 11] argued that linear optical response could detect clapping modes because the decay of an incident electromagnetic field at the boundary of a superconductor produces $\mathbf{q}$ components up to the inverse penetration depth. Bulk probes at $\mathbf{q}=0$ remain desirable., motivating techniques such as Raman [16, 17, 18, 19, 20, 21, 22, 23, 24, 25] or tunneling spectroscopy [26, 28, 29].
Previous studies have shown how linear electromagnetic coupling to superconducting collective modes can survive the long-wavelength limit in certain multi-band superconductors [38, 30, 14, 31, 32, 39], for example, or when the gap function breaks in-plane rotational symmetries [13]. An alternative is to apply a supercurrent [40, 41, 42, 43]: Cooper pairs then acquire a net velocity, coupling light to order-parameter amplitude fluctuations (i.e., the Higgs mode) even at $\mathbf{q}=0$.
In this Letter, we show how trigonal warping enables a similar effect for BaSch and clapping modes when pairing occurs around a single Fermi surface (see Fig. 1). These modes then become visible to any probe sensitive to the frequency-dependent superfluid stiffness, including longitudinal and Hall conductivity, polar Kerr effect [44], microwave reflectometry and related circuit-QED techniques [45, 46, 47, 48, 49, 50, 51]. Crucially, the optical couplings enabled by trigonal warping do not require any background field or supercurrent.
Figure 1: Trigonal warping of a single Fermi surface (a) corresponds to an inversion-even contribution $\tilde{\mathbf{v}}$ to the electron velocity (b). Zero-momentum Cooper pairs thus acquire a non-zero velocity, enabling linear optical response in the superconducting state.
@@
Rhombohedral four-, five-, and six-layer graphene (RnnG) may provide precisely the ingredients required for the optical couplings we discuss. Recent experiments [52, 53, 54] have revealed several superconducting phases that emerge from spin-and-valley-polarized normal states, each hosting a single $C_{3}$-symmetric Fermi surface. Theory suggests that some of these superconductors are chiral [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]. They might therefore host clapping modes (in addition to BaSch modes). We discuss the implications of our results for RnnG, and propose microwave impedance spectroscopy as an ideal probe of its superconducting collective modes.
@@
with attractive couplings $g_{1}\geq g_{2}>0$; the factor of $1/2$ is for convenience. (Including more channels is conceptually straightforward.) $\delta g=g_{1}-g_{2}$ quantifies the competition between pairing channels at the critical temperature; $\delta g=0$ means perfect competition.
@@
The gap equation arises as the saddle-point condition for the $\Delta^{(\ell)}$ fields after integrating out the fermions. Our analysis holds when inversion breaking is weak enough that $|\tilde{\xi}_{\mathbf{k}}|<E_{0,\mathbf{k}}\equiv\sqrt{\xi_{0,\mathbf{k}}^{2}+|\Delta_{\mathbf{k}}|^{2}}$ at all relevant momenta, where $\Delta_{\mathbf{k}}$ is the mean-field gap function (i.e., the saddle-point value of $\Delta_{\mathbf{k},q=0}$). For a nodeless state with strong pairing interactions, this condition can be satisfied self-consistently. In that case, surprisingly, $\tilde{\xi}$ has no impact on pairing and disappears from all calculations in the long-wavelength limit at $T\rightarrow 0$ [75], except for its contribution $\propto\tilde{\mathbf{v}}_{\mathbf{k}}\tau^{z}$ to the paramagnetic vertex (more on this below).
We assume that the $\ell=1$ channel condenses, producing the pairing gap
@@
$\Delta^{(1)}(x)=$ $e^{i\theta(x)}\left[\Delta_{\text{mf}}+h(x)\right],$ (7a)
@@
$h(x)$ and $\theta(x)\in\mathbb{R}$ are respectively the Higgs and Anderson-Bogoliubov-Goldstone modes. The interpretation of $a(x)$ and $b(x)\in\mathbb{R}$ depends on the pairing channels chosen. If $\chi^{(1)}_{\mathbf{k}}=\overline{\chi}^{(2)}_{\mathbf{k}}$, i.e., if the two pairing channels are time-reversed partners, then $a$ and $b$ are clapping modes. If instead $\chi^{(1)}_{\mathbf{k}}$ and $\chi^{(2)}_{\mathbf{k}}$ come from unrelated irreps, e.g., $p+ip$ and $f+if$ (or $s\text{-wave}$ and $d\text{-wave}$ in a spin-singlet model), then $a$ and $b$ would instead constitute BaSch modes. It will be useful to collect the fields $h,a$ and $b$ into $\Phi=\begin{pmatrix}h&a&b\end{pmatrix}^{T}$.
We now couple to electromagnetism through the standard minimal substitution. Electromagnetic linear response is controlled by the effective action at bilinear order in the gauge potential $A=(A_{0},\mathbf{A})$, so we drop all terms at third order or higher. We may remove the overall order-parameter phase $\theta(x)$ by a gauge transformation [76, 77, 78], which combines $A$ and the spacetime-gradient of $\theta$ into the gauge-invariant field $\mathcal{A}=(\mathcal{A}_{0},\boldsymbol{\mathcal{A}})\equiv(A_{0}+\frac{1}{2e}\partial_{\tau}\theta,\mathbf{A}-\frac{1}{2e}\gradient\theta)$. Note that $\theta$ will ultimately play no role in our calculation when we focus on $\mathbf{q}=0$ optical response [75]. We then reorganize the action for an expansion around mean field,
@@
Here $\tau^{0}$ and $\tau^{z}$ are the identity and third Pauli matrix in Nambu space. Equation (14) reveals an analogy to supercurrent-enabled optical coupling [40, 41, 79, 32, 43]: $\tilde{\mathbf{v}}_{\mathbf{k}}$ acts similarly to a background supercurrent at velocity $\mathbf{v}_{s}$, since both are inversion-even contributions to the electronic velocity which therefore contribute a $\tau^{z}$ to the paramagnetic vertex. Performing the Matsubara integral at $T\rightarrow 0$, we find
@@
where the 3-tuple is indexed by $\phi=h,a,b$. Equation (15) assumes only that $|\tilde{\xi}|<E_{0,\mathbf{k}}=\sqrt{\xi_{0,\mathbf{k}}^{2}+|\Delta_{\mathbf{k}}|^{2}}$ everywhere, and that $\Delta_{\mathbf{k}}=\Delta_{\text{mf}}\chi^{(1)}_{\mathbf{k}}$.
Unlike a constant background velocity, $\tilde{\mathbf{v}}_{\mathbf{k}}$ (which dictates the Cooper pair velocity) depends on direction. This is the key to enabling optical couplings to the $a$ and $b$ modes: the angle dependence of $\tilde{\mathbf{v}}_{\mathbf{k}}$ can compensate that of $\overline{\chi}^{(1)}_{\mathbf{k}}\chi^{(2)}_{\mathbf{k}}$, which a supercurrent cannot do. Conversely, while supercurrent optically activates the Higgs mode [40, 41, 42, 43], warping of the Fermi surface cannot since $\tilde{\mathbf{v}}_{\mathbf{k}}$ has a vanishing angular average (assuming that $\xi_{\mathbf{k}}$ respects any rotational symmetry $C_{n}$).
@@
$\tilde{\mathbf{v}}_{\mathbf{k}}=\partial_{\mathbf{k}}\tilde{\xi}_{\mathbf{k}}=\eta\mathbf{k}^{2}(\cos 2\varphi_{\mathbf{k}},-\sin 2\varphi_{\mathbf{k}}).$ (17)
Consider the couplings in Eq. (15). As explained above, the Higgs component vanishes since $\tilde{\mathbf{v}}_{\mathbf{k}}$ averages to zero over the Fermi surface, and the rest of the integrand is rotationally invariant. However, the clapping mode components include the factor $\overline{\chi}^{(1)}_{\mathbf{k}}\chi^{(2)}_{\mathbf{k}}=e^{-2i\varphi_{\mathbf{k}}}$ which balances the angular dependence of $\tilde{\mathbf{v}}_{\mathbf{k}}$. Those modes become optically active:
@@
In Fig. 2, we show the renormalized optical conductivity tensor
@@
along with the longitudinal stiffness, $K^{xx}_{\mathbf{q}=0}$, and the longitudinal inductance of a square sample, $\mathfrak{L}^{xx}$ (see [75]). The real part of the longitudinal conductivity consists of $\delta$-functions at DC ($\Omega=0$) and the clapping mode frequency $\Omega_{c}$, and a broad feature above the quasiparticle gap $2\Delta_{\text{mf}}$. Breaking the degeneracy between pairing channels ($g_{1}>g_{2}$) pushes the clapping mode up in energy from $\sqrt{2}\Delta_{\text{mf}}$. (Within our model, the $a$ and $b$ modes are degenerate for all values of $\delta g$, so they contribute a single pole; they may split when the gap function breaks rotational symmetry [13]). The Hall conductivity also exhibits distinctive features at $\Omega_{c}$ and $2\Delta_{\text{mf}}$. In clean superconductors, inversion symmetry typically prevents optical excitation of quasiparticles across the gap [80, 81, 82, 42] and enforces a vanishing Hall response [83, 84, 85]. By breaking inversion symmetry, trigonal warping activates both of these responses in our model.
[…]
@@
Application to RnnG—In rhombohedral graphene with four, five and six layers, experiments have reported superconductors that condense from spin-and-valley-polarized normal states [52, 53, 54], suggesting that pairing occurs within a single flavor. Fermion anticommutation then forces the gap function to be either chiral or nodal. Theoretical models often prefer a chiral solution as it admits a full gap, which increases the condensation energy relative to a nodal state [68, 59, 60, 56, 57, 58, 61, 62, 63, 64, 65, 66, 67]. Since the Fermi surface in each valley is trigonally warped, RnnG likely meets the criteria we have discussed for optical coupling to condensate modes.
In the SM we provide estimates for parameters appropriate for RnnG, used in Fig. 2. In particular, we predict a kinetic inductance on the order of tens of nH, at characteristic frequencies of a few GHz—natural scales for microwave reflectometry or circuit-QED inspired techniques. Recent measurements in few-layer NbSe2 [46, 50], MoTe2 [51] and moire graphene [48, 47, 49] have reported kinetic inductances on a similar scale.
@@
If a chiral superconductor spontaneously breaks $\mathcal{T}$ (i.e., if it arises from a $\mathcal{T}$-symmetric normal state), one generically expects a clapping mode comprising fluctuations in the time-reversed partner to the condensed channel. In superconductors descended from $\mathcal{T}$-broken normal states (like RnnG), the band-projected pairing interactions will also break $\mathcal{T}$, and attraction in the time-reversed channel could be much weaker than in the condensed channel. Another pairing channel (in a different irrep) could also have comparable attraction to the ground state [59, 73]. As a result, a BaSch mode might arise at lower energy than the clapping mode.
@@
Identifying the character of any collective mode observed in experiment could help gain insight into the underlying pairing interactions. The BaSch modes we considered change pair angular momentum by two units while preserving its direction, while the clapping mode reverses the pair angular momentum. Therefore, although the optical responses of these modes are identical in our model, we expect them to couple differently to an out-of-plane magnetic field. Future work will study this possibility within microscopic models.
@@
Our results extend the zoo of collective modes detectable in linear optical spectroscopy. We argued that valley-polarized superconductors in RnnG are a promising platform to test our ideas. Future work will consider more realistic models of RnnG and investigate how experimental tuning parameters—such as charge density, perpendicular displacement fields, and magnetic fields—affect their collective mode spectra.